OpenStudy (anonymous):

HELP, this determines my grade!! 1-) calculate the average rate of change for the function f(x)=x^4+3x^3-5x^2+2x-2 from x=-1 to x=1 2-) given the function f(x)=x^3+x^2-2x+1 what is the resulting function when f(x) is shifted to the left 1 unit 3-) given the function f(x)=-3x^3+9x^2-2x+3 what part of the function indicates that the left end starts at the top of the graph? 4-) graph the function f(x)=x^4-7x^3+12x^2+4x-12 and identifiy for which values of x the graph is increasing

OpenStudy (anonymous):

FIRST ONE IS \[\frac{f(1)-f(-1)}{1-(-1)}=\frac{-1-(-11)}{2}=\frac{10}{2}=5\]

OpenStudy (anonymous):

SECOND ONE IS \[f(x+1)=(x+1)^3+9(x+1)-2(x+1)+3=x^3+3 x^2+10 x+11\]

OpenStudy (anonymous):

Are you sure the first one isn't -5?